If x represents the first two consecutive even integers, express twice the sum of the two integers in terms of x and simplify? please help???
what i have come up with is 4x+4 ? I dont know if this is right?

Answer by rothauserc(4718)

In the general sense this problem can be written as
x, x+2 then twice the sum of the two consecutive even integers is
2 * (x + x+2) = 4x+4 = 4*(x+1)